• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name kuchina

    Log entry time 16:09:13 on May14,2007

    Entry number 202055

    keyword=Day shift summary

    During the shift we had no beam, so we took cosmics run.
    The problems we had:

    Low pressure alarm went off.

    Tried to move Right HRS to the new position. Right bogies didn't work.
    Scot Spiegel and Mark Stevens fixed it once. We moved spectrometer in wrong direction. Stopped the motion, could not start the bogies afterwards.

    Target was changed to BeO in preparation for the beam.
    Momentum for Right HRS is set for the new momentum.

    Right bogies problem and Left HRS momentum setup has been passed to the next shift.